# Enhancing Diffusion-Based Sampling with Molecular Collective Variables

### v1

@@ A.3 Molecular Isomerism
This appendix specifies how we operationalize chemical “sameness” and “difference” in terms of minimal geometric features, their coordination numbers at the atoms that define each stereochemical element, and the order parameter restraints we apply during sampling. We adopt the IUPAC taxonomy (McNaught & Wilkinson, 1997), which separates constitutional isomerism from stereoisomerism and further divides stereoisomerism into configurational and conformational classes.
@@
This taxonomy is paired with barrier and half-life delineations that mark fluxional, practically isolable, and configurationally stable regimes (Canfield et al., 2018).
@@
The polytope formalism in Canfield et al. (2018) provides a coordination-centric view of stereochemical change and shows how fundamental isomerization processes can be enumerated for each coordination number. In our context, this motivates representing each stereochemical element by the smallest stable geometric feature that distinguishes its alternatives, together with a numerical order parameter defined on those features. The order parameter then becomes the target of a light, topology-preserving restraint. Here, we exemplify this with two main types of restraints considered in this work, bond restraint and tetrahedral center chirality restraint. While not discussed here, other types of isomerism, such as axial, planar, and helical chirality, can be handled analogously.
@@ Appendix B Proofs and Notes
Proof. We follow Dama et al. (2014) for the bias decomposition and stochastic approximation framework, extending it with WT-ASBS-specific derivations and clearer notation.
@@
and $\zeta_{k+1}$ is a martingale difference with bounded second moment. The mini-batch hill update in Algorithm˜1 provides an unbiased estimator of the expectation $\mathbb{E}_{S\sim\bar{\nu}_{D}}$. From Step 1, $\epsilon_{k}\asymp\frac{1}{k}$, which satisfies the Robbins–Monro step condition $\sum_{k}\epsilon_{k}=\infty$ and $\sum_{k}\epsilon_{k}^{2}<\infty$ (Robbins & Monro, 1951). Standard stochastic approximation arguments then imply that the polygonal interpolation $\hat{D}(\tau_{k})=D_{k}$, $\hat{D}(\tau)=D_{k}+\frac{\tau-\tau_{k}}{\epsilon_{k}}(D_{k+1}-D_{k})$ for $\tau\in[\tau_{k},\tau_{k+1})$ is an asymptotic pseudo-trajectory of the mean-field ODE $\frac{\mathrm{d}}{\mathrm{d}\tau}D(\tau)=\Gamma[D(\tau)]$ (Kushner & Yin, 2003). Physically, this corresponds to the long-time limit equation in Dama et al. (2014), obtained by aggregating many small hill updates over a fixed $\tau$ window while the bias changes only slightly within that window.
@@ C.1 WT-ASBS with Replay Buffer
As proposed in ASBS (Liu et al., 2025), we warm-start the training process with the approximate bridge matching objective, which aligns the control model $u_{\theta}$ more closely with available data samples. In our setting, these samples are generated by short MD simulations that perform local sampling around a reference configuration $x_{\text{ref}}$. The bridge matching objective $\mathcal{L}_{\text{pretrain}}$ is defined as follows:
@@ C.2 Model Architecture
We build on PaiNN (Schütt et al., 2021), an $\mathrm{E}(3)$-equivariant graph neural network designed for molecular energy and force prediction, with minimal modifications: adding time-conditioning inputs, incorporating vector product layers (Schreiner et al., 2023) to break parity symmetry, restricting the model to $\mathrm{SE}(3)$-equivariance, and applying equivariant layer normalization (Liao et al., 2024) to improve training stability. Each atom is assigned a unique atomic embedding, as the reference configuration used to define the accessible component (Requirement˜1) already specifies atom indices. The model outputs a single vector per atom, which is scaled by $\sigma_{t}$ to produce the final control and corrector outputs.
@@ C.3 Base Process
We employ the base distribution in Karras et al. (2022) (EDM), which defines $\sigma_{t}$ through a power-schedule parameter $\rho$ and yields an analytically integrable variance between two different times $\kappa_{t|s}$:
@@
As in AS (Havens et al., 2025), we enforce translational invariance by projecting the SDE onto the subspace $\mathcal{X}^{\text{CoM}}=\{x\,|\,\sum_{i=1}^{n}x^{i}=0\}\subset\mathcal{X}$, where $x^{i}\in\mathbb{R}^{3}$ is the position of $i$-th atom. This is achieved through the projection operator $P$, which acts on any $y\in\mathbb{R}^{n\times 3}$ by subtracting the center of mass from each coordinate $y^{i}$ for $i\in[n]$.
@@ C.4 Hyperparameters
Hyperparameter
Alanine
dipeptide
Alanine tetrapeptide Nucleophilic substitution
Post-TS
bifurcation
@@
Initial Gaussian $h$ [eV] $1\times 10^{-4}$ 5 ×10^-4 $5\times 10^{-4}$ $2\times 10^{-4}$
@@
Model and training hyperparameters for each experiment are listed in Table˜1. For WT-ASBS Algorithm˜2, we use $L=100$ gradient updates per AM and CM epoch, and $\alpha_{\text{max}}=100.0$ to clip the energy gradient in the terminal cost. Models were trained with AdamW optimizer (Loshchilov & Hutter, 2017) using a weight decay of $10^{-3}$, and diffusion process was integrated with Euler–Maruyama method using 250 steps. For the first AM stage, we applied a 50-epoch learning rate warmup, after which bias deposition began. After completing $K$ training stages, we added a final AM stage without further bias deposition. Learning rates followed a cosine annealing schedule during training, applied independently to AM and CM optimization.
@@ Appendix D Experiment Details
For small peptides, we performed classical force field simulations with OpenMM 8.3.1 (Eastman et al., 2023), while reactions were modeled using UMA-S-1.1 uMLIP (Wood et al., 2025) as implemented in the fairchem-core 2.3.0 package (Shuaibi et al., 2025) with MD integrators in ASE 3.25.0 (Larsen et al., 2017). All MD-based enhanced sampling (WTMetaD) simulations were carried out with PLUMED 2.9.3 (Tribello et al., 2014; The PLUMED consortium, 2019).
@@ D.1 Alanine Dipeptide
The potential energy of the system was described by Amber ff96 force field (Kollman et al., 1997) with GBSA-OBC implicit solvation (Onufriev et al., 2004), as implemented in OpenMM (Eastman et al., 2023).
@@
We generated pretraining data using a 100 ps molecular dynamics simulation at 300 K with a 4 fs time step, enabled by hydrogen bond constraints and hydrogen mass repartitioning (HMR; Hopkins et al. (2015)) with hydrogen masses set to 1.5 amu. The system was equilibrated for 10 ps before sampling $10^{3}$ frames at 0.1 ps intervals. Simulations employed the Langevin middle integrator (Zhang et al., 2019) as implemented in OpenMM (Eastman et al., 2023), with a collision rate of 1 ps-1. In total, pretraining data generation required $2.75\times 10^{4}$ energy evaluations.
@@
We employed the test split of $10^{7}$ samples from Midgley et al. (2022), generated by replica exchange MD with 21 replicas spanning 300 K to 1300 K. Their dataset generation required $2.3\times 10^{10}$ energy evaluations.
@@
We performed WTMetaD simulations on the backbone torsion CVs $(\phi,\psi)$. The dynamics were propagated with Langevin middle integrator (Zhang et al., 2019) at 300 K with a 1 fs timestep and 1 ps-1 collision rate. Following a 10 ps equilibration at 300 K, Gaussian biases were deposited every 500 steps with widths of 0.2 rad for both variables. The bias update employed an initial Gaussian height of 1.2 kJ/mol, a bias factor of 8, and a grid covering $\pi,\pi]^{2}$ with $150\times 150$ bins. Each production run was 10 ns in length, and simulations were repeated four times to ensure statistical reliability, yielding $4.0\times 10^{7}$ energy evaluations.
@@ D.2 Alanine Tetrapeptide
The potential energy of the system was described by Amber ff99SB-ILDN force field (Lindorff-Larsen et al., 2010) with Amber99-OBC implicit solvation (Onufriev et al., 2004), as implemented in OpenMM (Eastman et al., 2023).
@@
We generated pretraining data using a 1 ns molecular dynamics simulation at 300 K with a 4 fs time step, enabled by hydrogen bond constraints and HMR with hydrogen masses set to 1.5 amu. The system was equilibrated for 10 ps before sampling $10^{3}$ frames at 1 ps intervals. Simulations employed the Langevin middle integrator (Zhang et al., 2019) as implemented in OpenMM (Eastman et al., 2023), with a collision rate of 1 ps-1. In total, pretraining data generation required $2.52\times 10^{5}$ energy evaluations.
@@
We generated unbiased alanine tetrapeptide conformations by first performing parallel tempering simulations to sample initial states, followed by long unbiased molecular dynamics runs for data collection. Parallel tempering simulations were carried out with OpenMM and OpenMMTools (Chodera et al., 2025) using MPI parallelization across 8 replicas, with temperatures geometrically spaced between 300 K and 600 K. Replicas were propagated with the generalized hybrid Monte Carlo (GHMC) integrator (Stoltz et al., 2010) using a 2 fs timestep (with hydrogen bond constraints), 1 ps-1 collision rate, and 1000 steps between exchange attempts (2 ps). Simulations were initialized from an energy-minimized structure, equilibrated for 200 ps, and then run for 100 ns of production with replica exchanges attempted every 2 ps.
From the 300 K replica, 64 independent unbiased MD simulations were seeded from randomly selected frames. Each simulation employed a Langevin middle integrator (Zhang et al., 2019) at 300 K with a 1 fs timestep and 1 ps-1 collision rate. After 100 ps of equilibration, production simulations were performed for 1.56 µs, yielding a total of 100 µs simulation time across all replicas. Trajectories were saved every 2 ps, producing $5\times 10^{7}$ samples in total. Reference data generation required approximately $1.01\times 10^{11}$ energy evaluations.
@@ D.3 Nucleophilic Substitution
The potential energy of the system was described by the UMA-S-1.1 (Wood et al., 2025) model with the DFT task label omol, a system charge of –1, and a singlet spin state (total spin 0).
@@ D.4 Post-Transition-State Bifurcation
The potential energy of the system was described by the UMA-S-1.1 (Wood et al., 2025) model with the DFT task label omol, a system charge of 0, and a singlet spin state (total spin 0).
@@
which approaches 1 when $d_{i}$ is much smaller than $d_{\text{ref}}$ and decays to 0 as $d_{i}$ increases (bond dissociation). To construct the CVs, we express the three relevant C–C distances (Figure˜4d) as normalized contacts using Equation˜55 with $d_{\text{ref}}=1.7$ Å, thereby capturing bond formation progress (Figure˜4e). We then define two CVs: $s_{1}=c_{1}+c_{2}+c_{3}$, which measures overall bond formation, and $s_{2}=c_{2}-c_{3}$, which distinguishes between products 4 and 5. A similar approach has been applied to enhanced sampling of mechanistically related reactions (Nam & Jung, 2023).
@@
We performed WTMetaD simulations using $(s_{1},s_{2})$ as the CVs. The dynamics were propagated with a Langevin thermostat at 300 K, a 0.5 fs timestep, and a friction coefficient of 10 ps-1. Initial velocities were assigned from a Maxwell–Boltzmann distribution, and the total momentum was removed. Following a 1 ps equilibration at 300 K, Gaussian biases were deposited every 250 steps with widths of 0.05 for both CVs. The biasing scheme used an initial Gaussian height of 0.02 eV, a bias factor of 20, and a grid spanning $0.0,3.0]\times[-1.0,1.0]$. Each production run was 0.8 ns in length, and simulations were repeated four times to ensure statistical reliability, yielding $6.4\times 10^{6}$ energy evaluations.
Appendix E Additional Results
E.1 Alanine dipeptide

### v2

@@ A.4 Molecular Isomerism
This appendix specifies how we operationalize chemical “sameness” and “difference” in terms of minimal geometric features, their coordination numbers at the atoms that define each stereochemical element, and the order parameter restraints we apply during sampling. We adopt the IUPAC taxonomy (mcnaught1997compendium), which separates constitutional isomerism from stereoisomerism and further divides stereoisomerism into configurational and conformational classes.
@@
This taxonomy is paired with barrier and half-life delineations that mark fluxional, practically isolable, and configurationally stable regimes (canfield2018new).
@@
The polytope formalism in canfield2018new provides a coordination-centric view of stereochemical change and shows how fundamental isomerization processes can be enumerated for each coordination number. In our context, this motivates representing each stereochemical element by the smallest stable geometric feature that distinguishes its alternatives, together with a numerical order parameter defined on those features. The order parameter then becomes the target of a light, topology-preserving restraint. Here, we exemplify this with two main types of restraints considered in this work, bond restraint and tetrahedral center chirality restraint. While not discussed here, other types of isomerism, such as axial, planar, and helical chirality, can be handled analogously.
@@ Appendix B Proofs and Notes
Proof. We follow dama2014well for the bias decomposition and stochastic approximation framework, extending it with WT-ASBS-specific derivations and clearer notation.
@@
and $\zeta_{k+1}$ is a martingale difference with bounded second moment. The mini-batch hill update in Algorithm˜1 provides an unbiased estimator of the expectation $\mathbb{E}_{S\sim\bar{\nu}_{D}}$. From Step 1, $\epsilon_{k}\asymp\frac{1}{k}$, which satisfies the Robbins–Monro step condition $\sum_{k}\epsilon_{k}=\infty$ and $\sum_{k}\epsilon_{k}^{2}<\infty$ (robbins1951stochastic). Standard stochastic approximation arguments then imply that the polygonal interpolation $\hat{D}(\tau_{k})=D_{k}$, $\hat{D}(\tau)=D_{k}+\frac{\tau-\tau_{k}}{\epsilon_{k}}(D_{k+1}-D_{k})$ for $\tau\in[\tau_{k},\tau_{k+1})$ is an asymptotic pseudo-trajectory of the mean-field ODE $\frac{\mathrm{d}}{\mathrm{d}\tau}D(\tau)=\Gamma[D(\tau)]$ (kushner2003stochastic). Physically, this corresponds to the long-time limit equation in dama2014well, obtained by aggregating many small hill updates over a fixed $\tau$ window while the bias changes only slightly within that window.
@@ C.1 WT-ASBS with Replay Buffer
As proposed in ASBS (liu2025adjoint), we warm-start the training process with the approximate bridge matching objective, which aligns the control model $u_{\theta}$ more closely with available data samples. In our setting, these samples are generated by short MD simulations that perform local sampling around a reference configuration $x_{\text{ref}}$. The bridge matching objective $\mathcal{L}_{\text{pretrain}}$ is defined as follows:
@@ C.2 Model Architecture
We build on PaiNN (schutt2021equivariant), an $\mathrm{E}(3)$-equivariant graph neural network designed for molecular energy and force prediction, with minimal modifications: adding time-conditioning inputs, incorporating vector product layers (schreiner2024implicit) to break parity symmetry, restricting the model to $\mathrm{SE}(3)$-equivariance, and applying equivariant layer normalization (liao2024equiformerv2) to improve training stability. Each atom is assigned a unique atomic embedding, as the reference configuration used to define the accessible component (Requirement˜1) already specifies atom indices. The model outputs a single vector per atom, which is scaled by $\sigma_{t}$ to produce the final control and corrector outputs.
@@ C.3 Base Process
We employ the base distribution in karras2022elucidating (EDM), which defines $\sigma_{t}$ through a power-schedule parameter $\rho$ and yields an analytically integrable variance between two different times $\kappa_{t|s}$:
@@
As in AS (havens2025adjoint), we enforce translational invariance by projecting the SDE onto the subspace $\mathcal{X}^{\text{CoM}}=\{x\,|\,\sum_{i=1}^{n}x^{i}=0\}\subset\mathcal{X}$, where $x^{i}\in\mathbb{R}^{3}$ is the position of $i$-th atom. This is achieved through the projection operator $P$, which acts on any $y\in\mathbb{R}^{n\times 3}$ by subtracting the center of mass from each coordinate $y^{i}$ for $i\in[n]$.
@@ C.4 Hyperparameters
Hyperparameter Alanine dipeptide Alanine tetrapeptide Nucleophilic substitution Post-TS bifurcation
@@
Initial Gaussian $h$ [eV] $5\times 10^{-4}$ 5 ×10^-4 $5\times 10^{-4}$ $2\times 10^{-4}$
@@
Model and training hyperparameters for each experiment are listed in Table˜1. For WT-ASBS Algorithm˜2, we use $L=100$ gradient updates per AM and CM epoch, and $\alpha_{\text{max}}=100.0$ to clip the energy gradient in the terminal cost. Models were trained with AdamW optimizer (loshchilov2017decoupled) using a weight decay of $10^{-3}$, and diffusion process was integrated with Euler–Maruyama method using 250 steps. For the first AM stage, we applied a 50-epoch learning rate warmup, after which bias deposition began. After completing $K$ training stages, we added a final AM stage without further bias deposition. Learning rates followed a cosine annealing schedule during training, applied independently to AM and CM optimization.
Several hyperparameters affecting training duration were selected based on the convergence of Algorithm˜2. The number of gradient updates $L$ was chosen to ensure that the controller parameters could track the outer-loop bias updates, as examined further in the ablation studies in Appendix˜E. The AM and CM epochs per stage ($M_{\text{adj}}$ and $M_{\text{crt}}$) were largely adopted from previous work (liu2025adjoint), and we found them sufficient for the AM and CM losses to stabilize within each training stage. The number of alternating training stages $K$ was set according to the convergence of the bias or PMF, assessed by whether newly deposited bias became negligible or whether the estimated free energy differences along the selected CVs had stabilized.
@@ Appendix D Experiment Details
For small peptides, we performed classical force field simulations with OpenMM 8.3.1 (eastman2023openmm), while reactions were modeled using UMA-S-1.1 uMLIP (wood2025uma) as implemented in the fairchem-core 2.3.0 package (fairchem) with MD integrators in ASE 3.25.0 (larsen2017atomic). All MD-based enhanced sampling (WTMetaD) simulations were carried out with PLUMED 2.9.3 (tribello2014plumed; plumed2019promoting).
@@ D.1 Alanine Dipeptide
The potential energy of the system was described by Amber ff96 force field (kollman1997development) with GBSA-OBC implicit solvation (onufriev2004exploring), as implemented in OpenMM (eastman2023openmm).
@@
We generated pretraining data using a 100 ps molecular dynamics simulation at 300 K with a 4 fs time step, enabled by hydrogen bond constraints and hydrogen mass repartitioning (HMR; hopkins2015long) with hydrogen masses set to 1.5 amu. The system was equilibrated for 10 ps before sampling $10^{3}$ frames at 0.1 ps intervals. Simulations employed the Langevin middle integrator (zhang2019unified) as implemented in OpenMM (eastman2023openmm), with a collision rate of 1 ps-1. In total, pretraining data generation required $2.75\times 10^{4}$ energy evaluations.
@@
We employed the test split of $10^{7}$ samples from midgley2022flow, generated by replica exchange MD with 21 replicas spanning 300 K to 1300 K. Their dataset generation required $2.3\times 10^{10}$ energy evaluations.
@@
We performed WTMetaD simulations on the backbone torsion CVs $(\phi,\psi)$. The dynamics were propagated with Langevin middle integrator (zhang2019unified) at 300 K with a 1 fs timestep and 1 ps-1 collision rate. Following a 10 ps equilibration at 300 K, Gaussian biases were deposited every 500 steps with widths of 0.2 rad for both variables. The bias update employed an initial Gaussian height of 1.2 kJ/mol, a bias factor of 8, and a grid covering $\pi,\pi]^{2}$ with $150\times 150$ bins. Each production run was 10 ns in length, and simulations were repeated four times to ensure statistical reliability, yielding $4.0\times 10^{7}$ energy evaluations.
@@ D.2 Alanine Tetrapeptide
The potential energy of the system was described by Amber ff99SB-ILDN force field (lindorff2010improved) with Amber99-OBC implicit solvation (onufriev2004exploring), as implemented in OpenMM (eastman2023openmm).
@@
We generated pretraining data using a 1 ns molecular dynamics simulation at 300 K with a 4 fs time step, enabled by hydrogen bond constraints and HMR with hydrogen masses set to 1.5 amu. The system was equilibrated for 10 ps before sampling $10^{3}$ frames at 1 ps intervals. Simulations employed the Langevin middle integrator (zhang2019unified) as implemented in OpenMM (eastman2023openmm), with a collision rate of 1 ps-1. In total, pretraining data generation required $2.52\times 10^{5}$ energy evaluations.
@@
We generated unbiased alanine tetrapeptide conformations by first performing parallel tempering simulations to sample initial states, followed by long unbiased molecular dynamics runs for data collection. Parallel tempering simulations were carried out with OpenMM and OpenMMTools (openmmtools) using MPI parallelization across 8 replicas, with temperatures geometrically spaced between 300 K and 600 K. Replicas were propagated with the generalized hybrid Monte Carlo (GHMC) integrator (stoltz2010free) using a 2 fs timestep (with hydrogen bond constraints), 1 ps-1 collision rate, and 1000 steps between exchange attempts (2 ps). Simulations were initialized from an energy-minimized structure, equilibrated for 200 ps, and then run for 100 ns of production with replica exchanges attempted every 2 ps.
From the 300 K replica, 64 independent unbiased MD simulations were seeded from randomly selected frames. Each simulation employed a Langevin middle integrator (zhang2019unified) at 300 K with a 1 fs timestep and 1 ps-1 collision rate. After 100 ps of equilibration, production simulations were performed for 1.56 µs, yielding a total of 100 µs simulation time across all replicas. Trajectories were saved every 2 ps, producing $5\times 10^{7}$ samples in total. Reference data generation required approximately $1.01\times 10^{11}$ energy evaluations.
@@ D.3 Nucleophilic Substitution
The potential energy of the system was described by the UMA-S-1.1 (wood2025uma) model with the DFT task label omol, a system charge of –1, and a singlet spin state (total spin 0).
@@ D.4 Post-Transition-State Bifurcation
The potential energy of the system was described by the UMA-S-1.1 (wood2025uma) model with the DFT task label omol, a system charge of 0, and a singlet spin state (total spin 0).
@@
which approaches 1 when $d_{i}$ is much smaller than $d_{\text{ref}}$ and decays to 0 as $d_{i}$ increases (bond dissociation). To construct the CVs, we express the three relevant C–C distances (Figure˜4d) as normalized contacts using Equation˜55 with $d_{\text{ref}}=1.7$ Å, thereby capturing bond formation progress (Figure˜4e). We then define two CVs: $s_{1}=c_{1}+c_{2}+c_{3}$, which measures overall bond formation, and $s_{2}=c_{2}-c_{3}$, which distinguishes between products 4 and 5. A similar approach has been applied to enhanced sampling of mechanistically related reactions (nam2023enhanced).
@@
We performed WTMetaD simulations using $(s_{1},s_{2})$ as the CVs. The dynamics were propagated with a Langevin thermostat at 300 K, a 0.5 fs timestep, and a friction coefficient of 10 ps-1. Initial velocities were assigned from a Maxwell–Boltzmann distribution, and the total momentum was removed. Following a 1 ps equilibration at 300 K, Gaussian biases were deposited every 250 steps with widths of 0.05 for both CVs. The biasing scheme used an initial Gaussian height of 0.02 eV, a bias factor of 30, and a grid spanning $0.0,3.0]\times[-1.0,1.0]$. Each production run was 0.8 ns in length, and simulations were repeated four times to ensure statistical reliability, yielding $6.4\times 10^{6}$ energy evaluations.
Appendix E Ablation Studies
E.1 Well-Tempered Bias Update
Here, we examine how the well-tempered bias hyperparameters (Gaussian height $h$, width $\sigma$, and bias factor $\gamma$) influence bias deposition, the intermediate sample distribution, and the convergence of free energy differences, using alanine dipeptide as a representative example.
Figure 7: Ablation of well-tempered bias updates on free energy convergence. The free energy difference $\Delta F$ between two states of alanine dipeptide is shown as a function of the number of energy evaluations during training. Each panel varies $h$, $\sigma$, and $\gamma$, and the default values used in the main text are indicated in gray background. $h$ and $\sigma$ are in units of eV and radians, respectively.
The convergence of $\Delta F$ shown in Figure˜7 across different well-tempered bias hyperparameters indicates that the method is robust over a broad range of settings. For the Gaussian height, values from $h=1.5\times 10^{-5}$ to $5\times 10^{-4}$ eV yield nearly identical converged results despite spanning almost a factor of 30, while only the largest value, $h=1.5\times 10^{-3}$ eV, shows a slight deviation. As seen in Figure˜8, increasing $h$ enhances early exploration of new modes and leads to a faster initial approach to the target $\Delta F$, but excessively large $h$ can impair final convergence. We note that, across different experiments, $h$ values on the order of $10^{-4}$ eV perform well. This is roughly two orders of magnitude smaller than the $h$ values commonly used in standard WTMetaD simulations, since multiple Gaussians are deposited simultaneously during WT-ASBS training.
Figure 8: Bias deposition and projected samples during training for different $h$ values. Each row shows, for selected $h$, how the deposited bias $V(\phi,\psi)$ (shifted to have a minimum of zero) evolves with increasing numbers of energy evaluations, along with the samples produced by the process at the corresponding training step, projected onto the $\phi$$\psi$ space. $h$ values are reported in units of eV.
For $\sigma$, larger values promote faster initial exploration by producing a broader bias early in training (Figure˜9), while their impact on $\Delta F$ convergence is comparatively modest (Figure˜7). For $\gamma$, we observe that a small value such as $\gamma=4$ leads to much slower convergence (Figure˜7) and reduced exploration (Figure˜10), as the effective temperature along the CVs is lower. For both $\sigma$ and $\gamma$, we observed that values commonly used in WTMetaD simulations provide good performance.
Figure 9: Bias deposition and projected samples during training for different $\sigma$ values. Each row shows, for selected $\sigma$, how the deposited bias $V(\phi,\psi)$ (shifted to have a minimum of zero) evolves with increasing numbers of energy evaluations, along with the samples produced by the process at the corresponding training step, projected onto the $\phi$$\psi$ space.
Figure 10: Bias deposition and projected samples during training for different $\gamma$ values. Each row shows, for selected $\gamma$, how the deposited bias $V(\phi,\psi)$ (shifted to have a minimum of zero) evolves with increasing numbers of energy evaluations, along with the samples produced by the process at the corresponding training step, projected onto the $\phi$$\psi$ space.
E.2 Training Mechanism
Now we perform analogous ablations on two training mechanisms central to the practical implementation of WT-ASBS: the replay buffer and data-based pretraining. Although several replay buffer hyperparameters, such as AM/CM buffer sizes, the number of buffer samples per epoch, and the number of gradient updates between buffer refreshes, can influence training dynamics, we use the number of AM gradient updates (Algorithm˜2, $L$ in line 8) as a representative control variable to assess the effect of the replay buffer. Because the replay buffer update is coupled to the bias update, increasing $L$ makes both updates less frequent.
Figure 11: Ablation of training mechanisms on free energy convergence. The free energy difference $\Delta F$ between two states of alanine dipeptide is shown as a function of the number of energy evaluations during training. Each panel varies the number of steps per epoch (replay buffer update frequency) and whether pretraining is used.
First, the $\Delta F$ convergence in Figure˜11 shows that the replay buffer has a substantial impact on training efficiency: using only a small number of AM steps per replay buffer update slows convergence significantly. The bias deposition patterns in Figure˜12 further indicate that overly frequent buffer and bias updates prevent the sampler parameters from keeping pace with the evolving bias, causing excessive bias to accumulate in the initial basin and leading to slow exploration of the next state. This illustrates that sufficient parameter updates between buffer updates help bridge the gap between the ideal two-time-scale WT-ASBS in Algorithm˜1, where ASBS training converges at each fixed bias, and the practical WT-ASBS in Algorithm˜2, where parameter and bias updates proceed concurrently. The replay mechanism must therefore provide enough time for parameter updates to keep up with bias deposition.
Figure 12: Bias deposition and projected samples during training for different number of AM steps per epoch. Each row shows, for selected number of steps per epoch, how the deposited bias $V(\phi,\psi)$ (shifted to have a minimum of zero) evolves with increasing numbers of energy evaluations, along with the samples produced by the process at the corresponding training step, projected onto the $\phi$$\psi$ space.
Next, Figure˜11 shows that $\Delta F$ convergence is substantially faster when training begins from localized pretraining. The sample distributions in Figure˜13 clarify this behavior: without pretraining, the model must identify low-energy modes from scratch, which requires a large number of energy gradient evaluations during the early stages of training. While the well-tempered bias update can converge to the correct PMF under mild conditions regardless of the initial sampling distribution, these results show that it is beneficial to exploit the initial structure of the chemical system and the efficiency of local, dynamics-based exploration.
Figure 13: Bias deposition and projected samples during training with or without pretraining. The upper row shows results without pretraining and the lower row shows results with pretraining. Each row illustrates how the deposited bias $V(\phi,\psi)$ (shifted to have a minimum of zero) evolves with increasing numbers of energy evaluations, along with the samples produced by the process at the corresponding training step, projected onto the $\phi$$\psi$ space.
Appendix F Additional Results
F.1 Alanine Dipeptide